OpenStudy (anonymous):

Find the slope of the line passing through the pair of points. (–4, 5) and (0, 2)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

slope is determined by y2-y1/x2-x1 (5-2)/(-4-0)=3/-4 or -3/4

OpenStudy (anonymous):

Find the slope of the line passing through the pair of points. (–2, –3) and (–2, –7)

OpenStudy (anonymous):

thank u i get it now

OpenStudy (anonymous):

I don't think so dalia_lam....a fraction of -4/0 has an undefined slope; it is a vertical line

OpenStudy (anonymous):

The answer isn't 0...do you see that?

OpenStudy (anonymous):

@DSS sorry, you're right. If the denominator is 0 the the slope is undefined.

OpenStudy (anonymous):



OpenStudy (anonymous):

Is the equation a direct variation? If it is, find the constant of variation. y = 2x

OpenStudy (anonymous):

????????

OpenStudy (anonymous):

yes y=kx k is the constant so y=2x 2 is the constant

OpenStudy (anonymous):

Is the equation a direct variation? If it is, find the constant of variation. y = x – 2

OpenStudy (anonymous):

Direct Variation: The statement " y varies directly as x ," means that when x increases, y increases by the same factor. In other words, y & x always have the same ratio: y/x= k where k is the constant of variation. We can also express the relationship between x and y as: y = kx where k is the constant of variation. y=x-2 is not a direct variation

OpenStudy (anonymous):

Is the equation a direct variation? If it is, find the constant of variation. Test Image

OpenStudy (anonymous):

y=x-2 is not a direct variation. It is a linear equation where the constant is -2

OpenStudy (anonymous):

Is the equation a direct variation? If it is, find the constant of variation.